Circular No. 6473

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1996 R2 Claes-Ingvar Lagerkvist, Uppsala Observatory, reports his discovery of

a comet on a Schmidt plate taken on Sept. 10 at the European Southern Observatory (ESO). As part of an Uppsala/DLR collaboration, on the following two nights U. Carsenty and S. Mottola obtained CCD observations with the ESO Bochum telescope that reveal a coma and anti-sunward tail.

1996 UT  α_{2000}  δ_{2000}  m_{1}
Sept.10.11667  22^{h}35^{m}48.^{s}5  −12^{o}40′30′′  17
11.18995  22 35 05.68  −12 43 55.2
11.19326  22 35 05.54  −12 43 55.6
12.08045  22 34 30.82  −12 46 31.1
12.08358  22 34 30.69  −12 46 31.5

A spectrum (range 400–800 nm) obtained by H. Pedersen with the Danish 1.54-m telescope at ESO revealed no obvious emission lines from the coma.

COMET P/1996 N2 (ELST-PIZARRO) R. H. McNaught, Anglo-Australian Observatory, has communicated

measurements of this object from a 75-min exposure by M. R. S. Hawkins with the U.K. Schmidt on 1985 Sept. 15, the faint, trailed image being asteroidal with V ∼ 19. These confirm B. G. Marsden’s suggested identification with 1979 OW_{7} (IAUC 6457); McNaught has also examined S. J. Bus’ 1979 U.K. Schmidt discovery plate, again confirming asteroidal character with V ∼ 18.5. Astrometric observations, orbital elements and an ephemeris are given on MPEC 1996-R07.

H. Böhnhardt, Institute for Astronomy and Astrophysics, University of Munich, reports: “Finson-Probstein calculations on the dust tail of P/1996 N2 indicate that, until late 1996, dust emitted before T + 40 to 60 days will be located on the sunward side of the nucleus. However, the discrimination between synchrones of different ages will be enhanced between mid-Sept. and mid-Dec.; from mid-Oct. to mid-Nov., a normal anti-sunward dust tail could develop if the dust-emission episode happened after T + 65 days. Earlier emission times may then be indicated by dust located over p.a. ∼ 300^{o}–252^{o}. Imaging during Oct.–Nov. could thus be very useful for constraining the time of the emission episode.”

Fully confirming the above, Z. Sekanina, Jet Propulsion Laboratory, notes that measures by W. Offutt on Sept. 9.15 UT give p.a. 254^{o}.03 ± 0^{o}.08, yielding an emission time of T + 42 days (formal error ± 4 days).

1996 September 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green